OpenStudy (anonymous):

which of the following equations are true? 1) -x +0+(- x) = 0 2) -x-(-x) =0 3) none of the above

OpenStudy (dayakar):

do u have any idea?

OpenStudy (anonymous):

1.) The +0+ doesn't have any value, so the equation simplifies to -x - x = 0, and when you combine the two -x's you get -2x = 0. 2.) -x - (-x) is the same thing as -x + x, so the equation simplifies to -x + x = 0. Since the positive x and the negative x cancel each other out, the equation becomes 0 = 0. So now we're left with: 1.) -2x = 0 2.) 0 = 0 I'll let you figure out which one is true ;)

OpenStudy (anonymous):

#2 is the correct answer

OpenStudy (anonymous):

yep